a sum of money is to be shared among three friends albert, ruth and christina in the ratio 3:7:10. if christina received $495 more than ruth determine the sum of money shared.
Answer by mananth(16946)

a sum of money is to be shared among three friends albert, ruth and christina in the ratio 3:7:10. if christina received $495 more than ruth determine the sum of money shared.
...
3:7:10
Total =20
Let total be $x
..
Ruth = 7/20*x
christina = 10/20 *x
(7x/20)+495 = 10x/20
(7x/20)+495=x/2
x/2-7x/20=495
(10x-7x)/20=495
3x=20*495
3x=9900
/3
x=3300
...
albert = 3/20 *3300= 495
Ruth = 7/20*3300=1155
Christina = 1/2 * 3300= 1650
CHECK
495+1155+1650=3300
...